# Adaptive variational quantum dynamics simulations with compressed circuits and fewer measurements

### v1

@@ Abstract
The adaptive variational quantum dynamics simulation (AVQDS) method performs real-time evolution of quantum states using automatically generated parameterized quantum circuits that often contain substantially fewer gates than Trotter circuits. Here we report an improved version of the method, which we call AVQDS(T), by porting the Tiling Efficient Trial Circuits with Rotations Implemented Simultaneously (TETRIS) technique. The algorithm adaptively adds layers of disjoint unitary gates to the ansatz circuit so as to keep the McLachlan distance, a measure of the accuracy of the variational dynamics, below a fixed threshold. We perform benchmark noiseless AVQDS(T) simulations of quench dynamics in local spin models demonstrating that the TETRIS technique significantly reduces the circuit depth and two-qubit gate count. We also show a method based on eigenvalue truncation to solve the linear equations of motion for the variational parameters with enhanced noise resilience. Finally, we propose a way to substantially alleviate the measurement overhead of AVQDS(T) while maintaining high accuracy by synergistically integrating quantum circuit calculations on quantum processing units with classical calculations using, e.g., tensor networks to evaluate the quantum geometric tensor. We showcase that this approach enables AVQDS(T) to deliver more accurate results than simulations using a fixed ansatz of comparable final depth for a significant time duration with fewer quantum resources.
@@ I Introduction
Quantum many-body systems have been a fruitful testbed for quantum computing Feynman (1982); Lloyd (1996); Abrams and Lloyd (1997, 1999); Somma et al. (2003); Aspuru-Guzik et al. (2005); Kassal et al. (2008); Georgescu et al. (2014); Wecker et al. (2015a); Lamm and Lawrence (2018); Smith et al. (2019); Chen et al. (2022); McArdle et al. (2020); Miessen et al. (2023), since they admit a natural mapping onto a qubit-based computing architecture Feynman (1982). In the near term, practical quantum algorithms need to be tailored for noisy intermediate-scale quantum (NISQ) hardware Preskill (2018) with limited qubits and coherence times. A typical example of such resource-efficient algorithms is the variational quantum eigensolver (VQE), which was introduced to solve static quantum problems such as finding the ground and excited states of a Hamiltonian by employing the variational principle Peruzzo et al. (2014); McClean et al. (2016); O’Malley et al. (2016); Kandala et al. (2017); Ryabinkin et al. (2018); Zhang et al. (2021). In VQE, parameterized trial states are prepared and measured on quantum processing units (QPUs), and a target cost function is calculated based on the measurement outcomes. The quantum circuits are then instructed by an optimization algorithm that runs on a classical computer on how to reach the optimal quantum state by updating the variational parameters. The variational ansatz plays a crucial role in bounding the accuracy of VQE. The problem-agnostic hardware efficient ansatz can in principle boost its expressibility by increasing the number of circuit layers, but practical applications are complicated by the presence of barren plateaus McClean et al. (2018); Larocca et al. (2024), where the cost function gradient vanishes exponentially with increasing number of qubits. The unitary coupled cluster ansatz with single and double excitations (UCCSD) is reasonably accurate to describe equilibrium ground states in quantum chemistry Barkoutsos et al. (2018), but the associated parameterized circuits can still be rather deep even for small molecules. One way of compressing the circuits while improving their ability to represent the ground state is to construct problem-specific ansätze by adaptively adding parameterized unitaries according to figures of merit such as the cost function gradient Grimsley et al. (2019) and state overlap Feniou et al. (2023). Examples include the adaptive derivative-assembled pseudo-trotter ansatz (ADAPT) approach and its variant known as qubit-ADAPT-VQE Grimsley et al. (2019); Tang et al. (2021); Feniou et al. (2023).
@@
Variational approaches with fixed ansätze have also been used to study the dynamics of quantum systems(VQDS) Yuan et al. (2019); Endo et al. (2020a); Cirstoiu et al. (2020); Commeau et al. (2020); Benedetti et al. (2021); Chen et al. (2019); Berthusen et al. (2022). Here, it becomes more challenging to construct the ansätze for dynamics simulations, because they are expected to faithfully represent the dynamical quantum state during the entire time evolution. For this reason, the adaptive variational quantum dynamics simulation (AVQDS) approach has been proposed Yao et al. (2021), which dynamically expands the variational ansatz to keep a figure of merit for the variational dynamics, known as the McLachlan distance $\mathcal{L}^{2}$, a measure of the fidelity of the time-evolved variational state, below a fixed threshold along the time-evolution path McLachlan (1964); Yuan et al. (2019). Compared with standard first-order Trotterized circuits for dynamics simulations of quantum spin models, it has been demonstrated that AVQDS can significantly reduce the circuit complexity as measured by the number of two-qubit entangling gates Yao et al. (2021); Mootz et al. (2024).
@@
In the original ADAPT-VQE and AVQDS methods, the parameterized unitary gates are added iteratively and one at a time when the ansatz is expanded. In general, each unitary only acts on a fraction of the qubits, leaving many of them idling. Adding the next unitary without taking the idle qubits into consideration can cause unnecessary deepening of the quantum circuit, making it suboptimal to run on QPUs. As a further improvement, Anastasiou et al. introduced the tiling efficient trial circuits with rotations implemented simultaneously (TETRIS) technique Anastasiou et al. (2022) to the original ADAPT-VQE method as an improved prescription for adding unitaries to the ansatz Tang et al. (2021). TETRIS begins like the standard ADAPT protocol by appending to the ansatz a unitary producing the largest energy gradient. The set of possible unitaries is characterized by a predefined operator pool consisting of the allowed generators of a unitary rotation. However, instead of proceeding with the VQE, the TETRIS approach keeps checking other unitaries from the pool with energy gradients ranked in descending order and adds the ones that act on disjoint sets of idling qubits until a full gate layer is obtained. In addition to the expected advantage of generating shallower ansatz circuits, the TETRIS technique was found to reduce the number of ansatz reoptimizations entailed in ADAPT-VQE calculations of small molecules, thereby speeding up convergence Anastasiou et al. (2022).
@@
In this work we harness the TETRIS technique for quantum dynamics simulations by incorporating it into the adaptive ansatz expansion subroutine of AVQDS, resulting in a modified version of the algorithm that we denote AVQDS(T). We apply AVQDS(T) to quench dynamics simulations of local spin models and show that this procedure greatly reduces the quantum circuit depth with negligible overhead. Meanwhile, we present the implementation of an approach based on eigenvalue truncation Hansen (1990); Gacon et al. (2023) to solve the linear equations of motion for propagating the variational parameters, and demonstrate that this method can be advantageous over standard techniques like Tikhonov regularization in the presence of noise.
@@
Furthermore, we propose an approach to synergistically integrate classical and quantum resources to address one of the main challenges of variational approaches like AVQDS. Specifically, AVQDS and related approaches all rely on measurements of the quantum geometric tensor (QGT) to propagate the variational parameters, necessitating the evaluation of a number of circuits that grows quadratically with the number of variational parameters Yuan et al. (2019). The pivotal idea is to capitalize on the observation that the quantum resource-hungry QGT can always be partitioned into an inner block, which can be evaluated efficiently and accurately using classical approaches like tensor networks (TNs), and the remaining components which are beyond the reach of TNs and therefore must be measured on QPUs. Finally, we showcase that with this implementation AVQDS(T) can deliver more accurate results than VQDS simulations using a comparable fixed ansatz over a significant time duration, while maintaining lower computational costs. The idea of hybrid QPU-TN evaluations of the QGT can be naturally extended to ground state preparation using imaginary time evolution McArdle et al. (2019); Gomes et al. (2021) and simulations of mixed states and open quantum systems Yuan et al. (2019); Chen et al. (2024). This allows the variational algorithms to maximally leverage the computational power of TNs and only resort to QPUs for the classically challenging part of the calculation, making them promising to achieve quantum utility in large-scale applications.
@@ II.1 Variational Quantum Dynamics Simulations
For completeness, we begin with a summary of the VQDS algorithm Yuan et al. (2019). The exact time evolution of the density matrix $\rho$ of a quantum state is governed by the von-Neumann equation
@@
which are derived by minimizing the McLachlan distance $\mathcal{L}^{2}$ between the exact and the variational time evolution McLachlan (1964):
@@ II.2 Adaptive Variational Quantum Dynamics Simulations (AVQDS)
where $\theta_{\mu}$ are variational parameters, $\hat{\mathcal{A}}_{\mu}$ are Hermitian operators selected from a pre-determined pool $\mathscr{P}=\{\hat{\mathcal{A}}_{i}\}_{i=1}^{N_{p}}$ made of $N_{p}$ Pauli strings, and |$\ket{\Psi_{0}}$ is a reference state. The unitaries $e^{-i\theta_{\mu}\hat{\mathcal{A}}_{\mu}}$ can be implemented on quantum devices with a combination of one-qubit rotation gates and two-qubit entangling gates Nielsen and Chuang (2011). In AVQDS, the ansatz is iteratively expanded during the evolution by adding new operators whenever the McLachlan distance exceeds a threshold value $\mathcal{L}^{2}_{\text{cut}}$ Yao et al. (2021). At each time step, one first solves Eq. (2) to obtain $\dot{\boldsymbol{\theta}}$ and the corresponding McLachlan distance $\mathcal{L}^{2}=2\left(\text{var}[\hat{\mathcal{H}}]-V^{\dagger}\dot{\boldsymbol{\theta}}\right)$, where $\text{var}[\hat{\mathcal{H}}]=\langle\Psi(\boldsymbol{\theta})|\,\hat{H}^{2}\,|\Psi(\boldsymbol{\theta})\rangle-\left[\langle\Psi(\boldsymbol{\theta})|\,\hat{\mathcal{H}}\,|\Psi(\boldsymbol{\theta})\rangle\right]^{2}$ is the energy variance of the variational state. If $\mathcal{L}^{2}<\mathcal{L}^{2}_{\text{cut}}$, indicating that the current ansatz still performs well, one proceeds to update the variational state by incrementing the variational parameters via the Euler method: $\Delta\boldsymbol{\theta}=\dot{\boldsymbol{\theta}}\Delta t$ Iserles (2009). Otherwise, an adaptive ansatz growth procedure is triggered to improve the ansatz expressivity, as illustrated in Fig. 1. The procedure starts by calculating a “score” for each generator $\hat{\mathcal{A}}_{i}$ in the pool, defined as the reduction in the McLachlan distance by appending the associated unitary to the current ansatz. Specifically, to evaluate the score of a unitary $e^{-i\theta\hat{\mathcal{A}}}$, it is first appended to the ansatz with $\theta$ set to zero. This leaves the state unchanged, but increases the dimension of the parameter space $N_{\boldsymbol{\theta}}$ by 1. Accordingly, $M$ is augmented with an extra row and column, and an extra element is appended to $V$ (note that the gradient with respect to $\theta$ generically does not vanish even if $\theta$ is set to zero). Because $M$ is symmetric, only the new column of $M$ and the new element of $V$ need to be evaluated. The McLachlan distance $\mathcal{L}^{2}$ for the updated ansatz can then be obtained, which is mathematically guaranteed to decrease or stay the same. With the score of each unitary evaluated, a set of unitaries is chosen (in a manner described below) to be appended to the ansatz so as to maximally reduce $\mathcal{L}^{2}$. This process is iterated until a satisfactory $\mathcal{L}^{2}<\mathcal{L}^{2}_{\textrm{cut}}$ is obtained.
@@
Three methods can be adopted for choosing new unitaries to expand the ansatz. Method 1, used in the original AVQDS Yao et al. (2021), chooses the unitary with the best score at each iteration, leaving most qubits idle. In contrast, we propose two new methods, 2 and 3, that take the circuit depth explicitly into account and represent two implementations of the TETRIS approach within AVQDS. In Method 2, a unitary with the best score is selected subject to the requirement that the unitary must act only on the idling quits of the current circuit layer. A new layer is started only if there are no idling qubits in the current layer or if all the unitaries acting on idling qubits fail to deliver any appreciable reduction in $\mathcal{L}^{2}$. This practice guarantees that the circuit grows layer by layer in a compact way. Because of the additional constraint, in principle Method 2 could require more iterations than Method 1 to to achieve $\mathcal{L}^{2}<\mathcal{L}_{\textrm{cut}}^{2}$. Method 3 avoids this potential increase in the number of iterations by adding $p\geq 1$ unitaries at each iteration. To do so, it ranks all unitaries according to their score and proceeds down the list, adding each unitary provided that it does not act on qubits that have already been acted upon by the previous unitaries. Thus, Method 3 introduces a new full circuit layer at each iteration, in contrast to a single unitary as in Methods 1 and 2. Due to the larger added number of variational degrees of freedom, Method 3 produces a greater reduction in the McLachlan distance $\mathcal{L}^{2}$ per iteration as compared to Methods 1 and 2. In practice, we find that Methods 2 and 3 produce circuits of similar depth. We therefore adopt Method 3 for all following AVQDS(T) simulations.

### v2

@@ Abstract
The adaptive variational quantum dynamics simulation (AVQDS) method performs real-time evolution of quantum states using automatically generated parameterized quantum circuits that often contain substantially fewer gates than Trotter circuits. Here we report an improved version of the method, which we call AVQDS(T), by porting the Tiling Efficient Trial Circuits with Rotations Implemented Simultaneously (TETRIS) technique. The algorithm adaptively adds layers of disjoint unitary gates to the ansatz circuit so as to keep the McLachlan distance, a measure of the accuracy of the variational dynamics, below a fixed threshold. We perform benchmark noiseless AVQDS(T) simulations of quench dynamics in local spin models, and compare with an alternative adaptive variational approach on quantum resource requirement. Quantum dynamical simulations implementing realistic noise channels are also reported. Finally, we propose a way to substantially alleviate the measurement overhead of AVQDS(T) while maintaining high accuracy by synergistically integrating quantum circuit calculations on quantum processing units with classical calculations using, e.g., tensor networks to evaluate the quantum geometric tensor. We showcase that this approach enables AVQDS(T) to deliver more accurate results than simulations using a fixed ansatz of comparable final depth for a significant time duration with fewer quantum resources.
@@ I Introduction
Quantum many-body systems have been a fruitful testbed for quantum computing [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], since they admit a natural mapping onto a qubit-based computing architecture [1]. In the near term, practical quantum algorithms need to be tailored for noisy intermediate-scale quantum (NISQ) hardware [15] with limited qubits and coherence times. A typical example of such resource-efficient algorithms is the variational quantum eigensolver (VQE), which was introduced to solve static quantum problems such as finding the ground and excited states of a Hamiltonian by employing the variational principle [16, 17, 18, 19, 20, 21]. In VQE, parameterized trial states are prepared and measured on quantum processing units (QPUs), and a target cost function is calculated based on the measurement outcomes. The quantum circuits are then instructed by an optimization algorithm that runs on a classical computer on how to reach the optimal quantum state by updating the variational parameters. The variational ansatz plays a crucial role in bounding the accuracy of VQE. The problem-agnostic hardware efficient ansatz can in principle boost its expressibility by increasing the number of circuit layers, but practical applications are complicated by the presence of barren plateaus [22, 23], where the cost function gradient vanishes exponentially with increasing number of qubits. The unitary coupled cluster ansatz with single and double excitations (UCCSD) is reasonably accurate to describe equilibrium ground states in quantum chemistry [24], but the associated parameterized circuits can still be rather deep even for small molecules. One way of compressing the circuits while improving their ability to represent the ground state is to construct problem-specific ansätze by adaptively adding parameterized unitaries according to figures of merit such as the cost function gradient [25] and state overlap [26]. Examples include the adaptive derivative-assembled pseudo-trotter ansatz (ADAPT) approach and its variant known as qubit-ADAPT-VQE [25, 27, 26].
@@
Variational approaches with fixed ansätze have also been used to study the dynamics of quantum systems(VQDS) [28, 29, 30, 31, 32, 33]. Here, it becomes more challenging to construct the ansätze for dynamics simulations, because they are expected to faithfully represent the dynamical quantum state during the entire time evolution. For this reason, the adaptive variational quantum dynamics simulation (AVQDS) approach has been proposed [34], which dynamically expands the variational ansatz to keep a figure of merit for the variational dynamics, known as the McLachlan distance $\mathcal{L}^{2}$, a measure of the fidelity of the time-evolved variational state, below a fixed threshold along the time-evolution path [35, 28]. Compared with standard first-order Trotterized circuits for dynamics simulations of quantum spin models, it has been demonstrated that AVQDS can significantly reduce the circuit complexity as measured by the number of two-qubit entangling gates [34, 36]. Alternative algorithms for compressing quantum circuits in dynamics simulations include projected variational quantum dynamics (pVQD) [37] and its adaptive variant [38], as well as variational fast forwarding (VFF) and the subspace approach [30, 39, 40, 41]. The adaptive pVQD algorithm iteratively constructs and optimizes a variational ansatz to maintain a high overlap with the quantum state evolved using one or more Trotter steps along the dynamical path. While both AVQDS and pVQD are suited for dynamics simulations with either fixed or time-dependent Hamiltonians, VFF and the subspace approach are tailored for sudden quench dynamics under time-independent Hamiltonians. The common underlying idea is to design variational circuits that approximately diagonalize the Hamiltonian evolution operator $e^{-i\hat{H}t}$ either across the entire Hilbert space or within a low-energy subspace, thereby simplifying the simulation.
@@
In the original ADAPT-VQE and AVQDS methods, the parameterized unitary gates are added iteratively and one at a time when the ansatz is expanded. In general, each unitary only acts on a fraction of the qubits, leaving many of them idling. Adding the next unitary without taking the idle qubits into consideration can cause unnecessary deepening of the quantum circuit, making it suboptimal to run on QPUs. As a further improvement, Anastasiou et al. introduced the tiling efficient trial circuits with rotations implemented simultaneously (TETRIS) technique [42] to the original ADAPT-VQE method as an improved prescription for adding unitaries to the ansatz [27]. TETRIS begins like the standard ADAPT protocol by appending to the ansatz a unitary producing the largest energy gradient. The set of possible unitaries is characterized by a predefined operator pool consisting of the allowed generators of a unitary rotation. However, instead of proceeding with the VQE, the TETRIS approach keeps checking other unitaries from the pool with energy gradients ranked in descending order and adds the ones that act on disjoint sets of idling qubits until a full gate layer is obtained. In addition to the expected advantage of generating shallower ansatz circuits, the TETRIS technique was found to reduce the number of ansatz reoptimizations entailed in ADAPT-VQE calculations of small molecules, thereby speeding up convergence [42].
@@
In this work we harness the TETRIS technique for quantum dynamics simulations by incorporating it into the adaptive ansatz expansion subroutine of AVQDS, resulting in a modified version of the algorithm that we denote AVQDS(T). We apply AVQDS(T) to quench dynamics simulations of local spin models and show that this procedure greatly reduces the quantum circuit depth with negligible overhead. Meanwhile, we present the implementation of an approach based on eigenvalue truncation [43, 44] to solve the linear equations of motion for propagating the variational parameters, and demonstrate that this method can be advantageous over standard techniques like Tikhonov regularization in the presence of noise.
@@
Furthermore, we propose an approach to synergistically integrate classical and quantum resources to address one of the main challenges of variational approaches like AVQDS. Specifically, AVQDS and related approaches all rely on measurements of the quantum geometric tensor (QGT) to propagate the variational parameters, necessitating the evaluation of a number of circuits that grows quadratically with the number of variational parameters [28]. The pivotal idea is to capitalize on the observation that the quantum resource-hungry QGT can always be partitioned into an inner block, which can be evaluated efficiently and accurately using classical approaches like tensor networks (TNs), and the remaining components which are beyond the reach of TNs and therefore must be measured on QPUs. Finally, we showcase that with this implementation AVQDS(T) can deliver more accurate results than VQDS simulations using a comparable fixed ansatz over a significant time duration, while maintaining lower computational costs. The idea of hybrid QPU-TN evaluations of the QGT can be naturally extended to ground state preparation using imaginary time evolution [45, 46] and simulations of mixed states and open quantum systems [28, 47]. This allows the variational algorithms to maximally leverage the computational power of TNs and only resort to QPUs for the classically challenging part of the calculation, making them promising to achieve quantum utility in large-scale applications.
@@ II.1 Variational Quantum Dynamics Simulations
For completeness, we begin with a summary of the VQDS algorithm [28]. The exact time evolution of the density matrix $\rho$ of a quantum state is governed by the von-Neumann equation
@@
which are derived by minimizing the McLachlan distance $\mathcal{L}^{2}$ between the exact and the variational time evolution [35]:
@@
In Appendix B, we discuss multiple approaches for solving the equations of motion given in Eq. 2. In the results presented below, we use the truncation method with $\varepsilon=10^{-6}$ and $\varepsilon=10^{-3}$ for noiseless and noisy simulations, respectively.
@@ II.2 Adaptive Variational Quantum Dynamics Simulations (AVQDS)
where $\theta_{\mu}$ are variational parameters, $\hat{\mathcal{A}}_{\mu}$ are Hermitian operators selected from a pre-determined pool $\mathscr{P}=\{\hat{\mathcal{A}}_{i}\}_{i=1}^{N_{p}}$ made of $N_{p}$ Pauli strings, and |$\ket{\Psi_{0}}$ is a reference state. Specifically, we adopt a pool comprising all the non-identity Pauli strings present in the Hamiltonian of the studied model. The unitaries $e^{-i\theta_{\mu}\hat{\mathcal{A}}_{\mu}}$ can be implemented on quantum devices with a combination of one-qubit rotation gates and two-qubit entangling gates [48]. In AVQDS, the ansatz is iteratively expanded during the evolution by adding new operators whenever the McLachlan distance exceeds a threshold value $\mathcal{L}^{2}_{\text{cut}}$ [34]. At each time step, one first solves Eq. (2) to obtain $\dot{\boldsymbol{\theta}}$ and the corresponding McLachlan distance $\mathcal{L}^{2}=2\left(\text{var}[\hat{\mathcal{H}}]-V^{\dagger}\dot{\boldsymbol{\theta}}\right)$, where $\text{var}[\hat{\mathcal{H}}]=\langle\Psi(\boldsymbol{\theta})|\,\hat{H}^{2}\,|\Psi(\boldsymbol{\theta})\rangle-\left[\langle\Psi(\boldsymbol{\theta})|\,\hat{\mathcal{H}}\,|\Psi(\boldsymbol{\theta})\rangle\right]^{2}$ is the energy variance of the variational state. If $\mathcal{L}^{2}<\mathcal{L}^{2}_{\text{cut}}$, indicating that the current ansatz still performs well, one proceeds to update the variational state by incrementing the variational parameters via the Euler method: $\Delta\boldsymbol{\theta}=\dot{\boldsymbol{\theta}}\Delta t$ [49]. Otherwise, an adaptive ansatz growth procedure is triggered to improve the ansatz expressivity, as illustrated in Fig. 1. The procedure starts by calculating a “score” for each generator $\hat{\mathcal{A}}_{i}$ in the pool, defined as the reduction in the McLachlan distance by appending the associated unitary to the current ansatz. Specifically, to evaluate the score of a unitary $e^{-i\theta\hat{\mathcal{A}}}$, it is first appended to the ansatz with $\theta$ set to zero. This leaves the state unchanged, but increases the dimension of the parameter space $N_{\boldsymbol{\theta}}$ by 1. Accordingly, $M$ is augmented with an extra row and column, and an extra element is appended to $V$ (note that the gradient with respect to $\theta$ generically does not vanish even if $\theta$ is set to zero). Because $M$ is symmetric, only the new column of $M$ and the new element of $V$ need to be evaluated. The McLachlan distance $\mathcal{L}^{2}$ for the updated ansatz can then be obtained, which is mathematically guaranteed to decrease or stay the same. With the score of each unitary evaluated, a set of unitaries is chosen (in a manner described below) to be appended to the ansatz so as to maximally reduce $\mathcal{L}^{2}$. This process is iterated until a satisfactory $\mathcal{L}^{2}<\mathcal{L}^{2}_{\textrm{cut}}$ is obtained.
@@
Three methods can be adopted for choosing new unitaries to expand the ansatz. Method 1, used in the original AVQDS [34], chooses the unitary with the best score at each iteration, leaving most qubits idle. In contrast, we propose two new methods, 2 and 3, that take the circuit depth explicitly into account and represent two implementations of the TETRIS approach within AVQDS. In Method 2, a unitary with the best score is selected subject to the requirement that the unitary must act only on the idling quits of the current circuit layer. A new layer is started only if there are no idling qubits in the current layer or if all the unitaries acting on idling qubits fail to deliver any appreciable reduction in $\mathcal{L}^{2}$. This practice guarantees that the circuit grows layer by layer in a compact way. Because of the additional constraint, in principle Method 2 could require more iterations than Method 1 to achieve $\mathcal{L}^{2}<\mathcal{L}_{\textrm{cut}}^{2}$. Method 3 avoids this potential increase in the number of iterations by adding $p\geq 1$ unitaries at each iteration. To do so, it ranks all unitaries according to their score and proceeds down the list, adding each unitary provided that it does not act on qubits that have already been acted upon by the previous unitaries. Thus, Method 3 introduces a new full circuit layer at each iteration, in contrast to a single unitary as in Methods 1 and 2. Due to the larger added number of variational degrees of freedom, Method 3 produces a greater reduction in the McLachlan distance $\mathcal{L}^{2}$ per iteration as compared to Methods 1 and 2. In practice, we find that Methods 2 and 3 produce circuits of similar depth. We therefore adopt Method 3 for all following AVQDS(T) simulations. In practice, we choose $\mathcal{L}_{\textrm{cut}}^{2}=10^{-3}$ for noiseless simulations and $\mathcal{L}_{\textrm{cut}}^{2}=10^{-1}$ for noisy simulations.